### v1

@@ 3 Chiti type comparison results
$\left.-\int_{R}^{r^{*}}f(r)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
@@
$\left.-f\left(r_{1}\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f\left(r_{1}\right)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
@@
We now prove that equality holds only when the domain $\Omega$ is a ball. Since $\left|B_{R}\right|<|\Omega|$, it follows that $R<r^{*}$. We claim that
@@
and the reverse inequality holds if $f$ is strictly decreasing on $[0,R]$, where $f$ is extended to $R,r^{*}]$ as the constant $f(R)$ with $f(R)>0$. Indeed, if $z(r)\geq u_{1}^{\star}(r)$ in $\left[0,R\right]$, there are two case: either $z(r)\equiv u_{1}^{\star}(r)$ in $\left[0,R\right]$ or $z(r)\geq u_{1}^{\star}(r)$ and $z(r)\not\equiv u_{1}^{\star}(r)$ in $\left[0,R\right]$. In fact, the former case is impossible. Otherwise, we find that
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$
@@
$=$ $NC_{N}\left(\int_{0}^{R}f(r)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-f(R)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
@@
$<\,$ $NC_{N}\left(\int_{0}^{R}f(R)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}f(R)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f(R)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
@@
$\left.-f\left(R\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
Since $f$ is strictly increasing, $r_{1}<R<r^{*}$ and $u_{1}^{\star}>0$, it follows that
$f\left(R\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r>f\left(r_{1}\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r.$
@@
$\left.-f\left(r_{1}\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$

### v2

@@ 3 Chiti type comparison results
$\left.-\int_{R}^{R_{*}}f(r)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
@@
$\left.-f\left(r_{1}\right)\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f\left(r_{1}\right)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
@@
We now prove that equality holds only when the domain $\Omega$ is a ball. Since $\left|B_{R}\right|<|\Omega|$, it follows that $R<R_{*}$. We claim that
@@
and the reverse inequality holds if $f$ is strictly decreasing on $[0,R]$, where $f$ is extended to $R,R_{*}]$ as the constant $f(R)$ with $f(R)>0$. Indeed, if $z(r)\geq u_{1}^{\star}(r)$ in $\left[0,R\right]$, there are two case: either $z(r)\equiv u_{1}^{\star}(r)$ in $\left[0,R\right]$ or $z(r)\geq u_{1}^{\star}(r)$ and $z(r)\not\equiv u_{1}^{\star}(r)$ in $\left[0,R\right]$. In fact, the former case is impossible. Otherwise, we find that
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{\left|\Omega\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$
@@
$=$ $NC_{N}\left(\int_{0}^{R}f(r)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-f(R)\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
@@
$<\,$ $NC_{N}\left(\int_{0}^{R}f(R)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{R_{*}}f(R)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f(R)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
@@
$\left.-f\left(R\right)\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
Since $f$ is strictly increasing, $r_{1}<R<R_{*}$ and $u_{1}^{\star}>0$, it follows that
$f\left(R\right)\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r>f\left(r_{1}\right)\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r.$
@@
$\left.-f\left(r_{1}\right)\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
@@
5 Proof of Theorem 1.2
A standard scaling argument shows that $\alpha^{2}/\widetilde{R}^{2}$ is the first eigenvalue of the corresponding problem on the ball $B_{\widetilde{R}}$
$\left\{\begin{array}[]{ll}-\Delta z=\mu z&\text{in}\,\,\,B_{\widetilde{R}},\\ \partial_{\texttt{n}}z+\frac{\sigma}{\widetilde{R}}z=0&\text{on}\,\,\,\partial B_{\widetilde{R}}\end{array}\right$.
with the corresponding eigenfunction
$z(r)=cr^{-\nu}J_{\nu}\left(\frac{\alpha}{\widetilde{R}}r\right),$
where $c$ is a positive constant. Since $\alpha=k_{\nu,1}$, we derive that $z\left(\widetilde{R}\right)>0$.
Similar to Lemma 3.2, we have the following relation between the measures of $B_{\widetilde{R}}$ and $\Omega$.
Lemma 5.2. Assume that $\widetilde{R}\leq R$. It holds that $\left|B_{\widetilde{R}}\right|\leq|\Omega|$, with equality if and only if $\Omega=B_{\widetilde{R}}$.
Proof. The condition $\widetilde{R}\leq R$, combined with Lemma 3.2, shows that
$\left|B_{\widetilde{R}}\right|\leq\left|B_{{R}}\right|\leq|\Omega|.$
When $\Omega$ is a ball, we find that $\widetilde{R}=R_{*}=R$, and thus equality holds in the inequality above. On the other hand, if $\Omega$ is not a ball, Lemma 3.2 implies that $\left|B_{{R}}\right|<|\Omega|$, from which it follows that $\left|B_{\widetilde{R}}\right|<|\Omega|$. Therefore, we conclude that $\left|B_{\widetilde{R}}\right|\leq|\Omega|$, with equality if and only if $\Omega=B_{\widetilde{R}}$.∎
By Lemma 5.2 and Lemma 3.2, if $\Omega$ is a ball, one sees that $\widetilde{R}=R_{*}=R$. In this case, the conclusion of Theorem 1.2 holds trivially (by letting $R_{*}=R$) for all $\sigma>0$. We may therefore assume throughout the remainder of this section that $\left|B_{\widetilde{R}}\right|<|\Omega|$ and $\widetilde{R}\leq R$.
[…]
$\left(u_{1}^{*}\right)^{\prime}(s)\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{2/N-2}\int_{0}^{s}u_{1}^{*}(t)\,\text{d}t$
for almost every $s\in\left[0,s_{*}\right]$. Then, repeating the arguments of Lemmas 3.3–3.4, Theorem 3.1 and Theorem 1.1 by replacing $R$ with $\widetilde{R}$, we can obtain that
$\mu_{2}-\mu_{1}\leq\frac{1}{\widetilde{R}^{2}}\left(\beta^{2}-\alpha^{2}\right),$
which implies that
$\frac{\mu_{2}}{\mu_{1}}<\frac{k_{\nu+1,1}^{2}-k_{\nu,1}^{2}}{\widetilde{R}^{2}\mu_{1}}+1,$
which is the desired inequality.∎